A painter can paint a house in 3 days. When his helper works with him, both men can paint the house in 2 days. How long will it take the helper, working alone, to paint the house?
Answer by rfadrogane(214)

A painter can paint a house in 3 days. When his helper works with him, both men can paint the house in 2 days. How long will it take the helper, working alone, to paint the house?
Sol'n:
let x: the number of days the helper work alone
rate x time = 1
1/3+1/x=1/2
1/x=1/6
x = 6 days ----answer